trying to just double check the rest of my daughters review
foiling expressions

(x+3)(x^2+3x-2) I came up with 3x^4+3-2
It does not sound right but I tried the answer my daughter has is 3x^2+3-2

Answer by stanbon(48526)

(x+3)(x^2+3x-2)
=x(x^2+3x-2)+3(x^2+3x-2)
=x^3+3x^2-2x+3x^2+9x-6
=x^3+6x^2+7x-6
Good luck with your work.
Cheers,
Stan H.